# Heapr: Hessian-based Efficient atomic expert Pruning in Output Space

### v1

@@ Appendix A Derivation of the importance of atomic expert
Consider the negative log-likelihood loss $\ell$, whose per-sample gradient with respect to the parameters $\Theta$ can be written as
$\nabla_{\Theta}\ell=\bm{J}_{k}^{\top}\mathbf{g}_{\mathbf{e}_{k}},$ (17)
@@
where $\bm{J}_{k}\in\mathbb{R}^{d\times P}$ is the Jacobian of the atomic expert output $\mathbf{e}_{k}$ with respect to its parameters $\Theta$, and $\mathbf{g}_{\mathbf{e}_{k}}\in\mathbb{R}^{d}$ is the gradient of the loss with respect to $\mathbf{e}_{k}$. By definition, the Fisher Information Matrix is
$\bar{\bm{H}}=\bm{F}=\mathbb{E}\big[(\nabla_{\Theta}\ell)(\nabla_{\Theta}\ell)^{\top}\big].$ (18)
After model convergence, the Jacobian $\bm{J}_{k}$ can be treated as independent (Martens & Grosse, 2015), substituting the expression of $\nabla_{\Theta}\ell$ gives
$\bm{F}=\bm{J}_{k}^{\top}\,\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\bm{J}_{k}.$ (19)
Returning to the quadratic optimization problem in the OBS framework:
$\min_{\delta\Theta}\ \tfrac{1}{2}\,\delta\Theta^{\top}\bm{F}\,\delta\Theta\quad\text{s.t.}\quad\bm{J}_{k}\delta\Theta+\mathbf{e}_{k}=0,$ (20)
we define the auxiliary variable $\mathbf{u}=\bm{J}_{k}\delta\Theta$. The constraint becomes $\mathbf{u}+\mathbf{e}_{k}=0$, i.e., $\mathbf{u}=-\mathbf{e}_{k}$, and the objective reduces to
@@
$\min_{\mathbf{u}}\ \tfrac{1}{2}\,\mathbf{u}^{\top}\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{u}\quad\text{s.t.}\quad\mathbf{u}+\mathbf{e}_{k}=0.$ (21)
Plugging in the constraint yields the optimal cost:
$\Delta\ell=\tfrac{1}{2}\,\mathbf{e}_{k}^{\top}\,\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{e}_{k}.$ (22)
We therefore define the importance of the $k$-th atomic expert as
$s\;=\;\tfrac{1}{2}\,\mathbf{e}_{k}^{\top}\,\mathbb{E}\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{e}_{k}$ (23)
which is a scalar since $\mathbf{e}_{k}\in\mathbb{R}^{d}$ and $\mathbb{E}[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}]\in\mathbb{R}^{d\times d}$. This formalizes equation 13 in the main text: the smaller the value of $s$, the less impact the $k$-th atomic expert has on the overall model loss, making it a better candidate for pruning.

### v2

@@ Acknowledgments
  • Section A: Derivation of the importance of atomic expert.
  • Section B: Detail Analysis of Main Results.
  • Section C: Analysis of Runtime Speedup and Memory Usage.
  • Section D: Compression Rate Analysis under Global Pruning.
  • Section E: Use of LLM.
  • Section F: Reproducibility Statement.
  • Section G: Ethics statement.
@@ Appendix A Derivation of the Importance of Atomic Expert
Let $\Theta$ denote the model parameters of dimension $P$. Let $\mathbf{e}_{k}(\mathbf{x})\in\mathbb{R}^{d}$ be the output of the $k$-th atomic expert for a given input $\mathbf{x}$. We define the Jacobian of the atomic expert output with respect to the parameters as $\bm{J}_{k}(\mathbf{x})=\frac{\partial\mathbf{e}_{k}(\mathbf{x})}{\partial\Theta}\in\mathbb{R}^{d\times P}$. Let $\mathbf{g}(y)=\nabla_{\mathbf{e}_{k}}\log P(y|\mathbf{x},\Theta)\in\mathbb{R}^{d}$ be the gradient of the log-likelihood with respect to the expert output.
Under the negative log-likelihood loss, the Hessian matrix $H$ can be approximated by the Fisher Information Matrix $\bm{F}$:
@@
$\bm{H}\approx\bm{F}=\mathbb{E}_{\mathbf{x}\sim P_{\text{data}}}\!\left[\mathbb{E}_{y\sim P_{\text{model}}(y|\mathbf{x})}\!\left[\nabla_{\Theta}\log P(y|\mathbf{x},\Theta)\nabla_{\Theta}\log P(y|\mathbf{x},\Theta)^{\top}\right]\right].$ (17)
Using the chain rule, the gradient of the log-likelihood with respect to the parameters can be decomposed as:
$\nabla_{\Theta}\log P(y|\mathbf{x},\Theta)=\bm{J}_{k}(\mathbf{x})^{\top}\mathbf{g}(y).$ (18)
Substituting this into the expression for $\bm{F}$:
$\bm{F}=\mathbb{E}_{\mathbf{x}}\!\left[\mathbb{E}_{y|\mathbf{x}}\!\left[\bm{J}_{k}(\mathbf{x})^{\top}\mathbf{g}(y)\mathbf{g}(y)^{\top}\bm{J}_{k}(\mathbf{x})\right]\right].$ (19)
Since $\bm{J}_{k}(\mathbf{x})$ depends only on $\mathbf{x}$ and is constant with respect to the inner expectation over $y$, we can extract it:
$\bm{F}=\mathbb{E}_{\mathbf{x}}\!\left[\bm{J}_{k}(\mathbf{x})^{\top}\left(\mathbb{E}_{y|\mathbf{x}}\!\left[\mathbf{g}(y)\mathbf{g}(y)^{\top}\right]\right)\bm{J}_{k}(\mathbf{x})\right].$ (20)
We define $\bm{G}_{k}(\mathbf{x})\triangleq\mathbb{E}_{y|\mathbf{x}}\!\left[\mathbf{g}(y)\mathbf{g}(y)^{\top}\right]$. Then, the parameter-space FIM becomes:
$\bm{F}=\mathbb{E}_{\mathbf{x}}\!\left[\bm{J}_{k}(\mathbf{x})^{\top}\bm{G}_{k}(\mathbf{x})\bm{J}_{k}(\mathbf{x})\right].$ (21)
We aim to find a parameter perturbation $\delta\Theta$ that minimizes the change in the overall loss while canceling the output of the $k$-th atomic expert (i.e. $\bm{J}_{k}(\mathbf{x})\delta\Theta+\mathbf{e}_{k}(\mathbf{x})=0$). The optimization problem is formulated as:
$\min_{\delta\Theta}\ \tfrac{1}{2}\,\delta\Theta^{\top}\bm{F}\,\delta\Theta\qquad\text{s.t.}\qquad\bm{J}_{k}(\mathbf{x})\,\delta\Theta+\mathbf{e}_{k}(\mathbf{x})=0.$ (22)
Substituting the derived form of $\bm{F}$:
$\min_{\delta\Theta}\ \tfrac{1}{2}\,\delta\Theta^{\top}\mathbb{E}_{\mathbf{x}}\!\left[\bm{J}_{k}(\mathbf{x})^{\top}\bm{G}_{k}(\mathbf{x})\bm{J}_{k}(\mathbf{x})\right]\delta\Theta.$ (23)
Since $\delta\Theta$ is a global parameter update independent of the specific sample $\mathbf{x}$, we can move it inside the expectation:
$\min_{\delta\Theta}\ \tfrac{1}{2}\,\mathbb{E}_{\mathbf{x}}\!\left[(\bm{J}_{k}(\mathbf{x})\delta\Theta)^{\top}\bm{G}_{k}(\mathbf{x})(\bm{J}_{k}(\mathbf{x})\delta\Theta)\right].$ (24)
Using the constraint $\bm{J}_{k}(\mathbf{x})\delta\Theta=-\mathbf{e}_{k}(\mathbf{x})$, we substitute this into the objective function:
$S_{k}=\tfrac{1}{2}\,\mathbb{E}_{\mathbf{x}}\!\left[\mathbf{e}_{k}(\mathbf{x})^{\top}\bm{G}_{k}(\mathbf{x})\mathbf{e}_{k}(\mathbf{x})\right].$ (25)
Appendix B Detail Analysis of Main Results.
Calibration Set Sampling Strategy.
[…]
Details of Baseline Experiments.
For DeepSeekMoE-16B-Base, the results for NAEE, MoE-I2, and $D^{2}$-MoE are taken from the paper (Gu et al., 2025), while MoE-SVD results are sourced from its paper (Li et al., 2025b). For Qwen1.5-MoE-A2.7B-Chat, all results are from the paper (Li et al., 2025a); any missing results with available official open-source code were reproduced by us. For Qwen3-30B-A3B, the results for HC-SMoE and Sub-MoE are from the paper (Li et al., 2025a). For Qwen2-57B-A14B, the results for NAEE, MoE-I2, and $D^{2}$-MoE are taken from the paper (Gu et al., 2025). Table 4 shows the calibration dataset size for various methods.
Table 4: Calibration set sizes for different methods (2048 sqlen).
Method NAEE $D^{2}$-MoE Sub-MoE HEAPr
Calibration Set Size 128 512 128 128
Appendix C Analysis of Runtime Speedup and Memory Usage
Table 5 summarizes the computational cost and performance of HEAPr compared with competitive baseline methods (NAEE and D2-MoE) on two representative MoE models: DeepSeekMoE-16B-Base and Qwen2-57B-A14B. The table reports the number of calibration samples used for pruning, the theoretical FLOPs (TFLOPs) required for pruning, GPU time cost, peak memory usage.
Table 5: Comparison of computational cost between HEAPr and baseline pruning methods.
Method Samples TFLOPs GPU Time Cost Memory
DeepSeekMoE-16B-Base
NAEE 128 11 2 min 27GB
D2-MoE 512 227 30 min 53GB
HEAPr 128 44 6 min 44GB
Qwen2-57B-A14B
NAEE 128 32 8 min 60GB
D2-MoE 512 1205 90 min 127GB
HEAPr 128 123 20 min 91GB
Appendix D Compression Rate Analysis under Global Pruning
In this section, we analyze the compression rates across different layers when applying a $25\%$ and $50\%$ global pruning strategy based on the global ranking of atomic experts’ importance. As shown in Figure 5 and  6, the compression rate is initially high in the early layers, suggesting that the experts in these layers are less important and can be pruned with minimal impact on the model’s performance. As we move deeper into the network, the compression rate decreases, indicating that the experts in these layers are more important to the model’s performance. Interestingly, after a certain point, the compression rate starts to increase again in the deepest layers, suggesting that some experts in these layers become redundant, allowing for further pruning without significant loss of model performance. This non-monotonic behavior highlights the varying importance of experts across layers in MoE-based models.
Figure 5: Compression ratios across different layers under 25% global pruning for Qwen1.5-MoE-A2.7B-Chat, DeepSeekMoE-16b-Base, and Qwen3-30B-A3B.
Figure 6: Compression ratios across different layers under 50% global pruning for Qwen1.5-MoE-A2.7B-Chat, DeepSeekMoE-16b-Base, and Qwen3-30B-A3B.
Appendix E Use of LLMs
In this work, Large Language Models (LLMs) were primarily utilized for tasks such as text refinement, offering writing suggestions, and improving the overall structure and clarity of the manuscript. It is important to note that LLMs did not contribute to the ideation or development of the methodology section. The authors guarantee that all LLM-generated content was thoroughly reviewed and edited to ensure its accuracy and coherence.
Appendix F Reproducibility Statement
To ensure reproducibility, we have made the code and checkpoints obtained in our computational environment available at https://github.com/LLIKKE/HEAPr. While we have taken every effort to ensure consistency, results may exhibit slight variations due to the random selection of calibration sets, as well as potential version differences in libraries such as transformers and LM-Evaluation-Harness. These fluctuations are expected and considered acceptable.
Appendix G Ethics Statement
This work adheres to ethical guidelines in conducting research and reporting results. We have used publicly available datasets and models, ensuring that our methods comply with their respective terms of use. The research itself aims to enhance existing technologies and does not introduce any ethical concerns. No personal or sensitive data was used in this study, and the methods employed do not raise any known ethical issues.